OpenStudy (anonymous):

Write the equation of the line which passes through (3, –4) and is perpendicular to the line with equation 4y – x = 1.

OpenStudy (anonymous):

Equation is in the form y=mx+b. You know you want it to be perpendicular, which means its slope m should be the negative inverse of the equation given... 4y-x = 1 4y = x+1 y = x/4 + 1/4 So the slope of the equation given is 1/4. This means the slope of the equation we want should be -4: the negative inverse. Substituting in the point (3, -4) we have yields: y = mx+b (-4) = (-4)(3)+b -4 = -12 + b b = 8 Therefore the final equation we get is: y = -4x + 8 Hope this helps! 8D

OpenStudy (anonymous):

It does, thanks